# Nonuniqueness of Leray–Hopf solutions to the unforced incompressible 3D Navier–Stokes Equation

### v1

@@ 6 Finite Rank Approximation
$Q_{s}=\frac{\nabla\overline{U}+\nabla^{T}\overline{U}}{2},\quad Q_{s}+\eta_{1}I=Q_{p}-Q_{m},\quad Q_{p}\succeq 0,\quad Q_{m}\succeq 0.$
Here, $Q_{m}$ is the negative part of $Q_{s}+\eta_{1}I$. In previous sections, we choose $Q=Q_{m}$. We truncate the domain $\Omega=[0,R_{0}]\times\left[0,\frac{\pi}{2}\right]$ with $R_{0}=32$ and verify that outside $\Omega$, we have $\lambda_{\max}(Q_{s})<0.191$. Thus, we set $\eta_{1}=0.191$ so that the support of $Q_{m}$ is fully contained in $\Omega$.
Although it is natural to take $Q=Q_{m}$, this choice has the drawback that $Q$ is not smooth, which complicates the future error analysis of numerical integrals. Fortunately, we only need $Q$ to satisfy
@@ 6.1 Construction of $Q$
A naïve approach would be to apply the discrete Fourier transform to $Q_{m}$ and truncate the series. However, this does not ensure $Q_{s}+\eta_{1}I+Q\succeq 0$. To address this, we first select a slightly smaller $\eta_{0}=0.19$ and compute the corresponding $Q_{m}$. Then we reconstruct $Q$ using the same idea. As long as $L,M$ are sufficiently large, we obtain (6.1).
This method, however, typically requires very large $L,M$, which is undesirable. To improve efficiency, we fix $L=150,M=300$ and initialize $Q_{0}=0$ with $\eta_{0}=0.19$. For every entry $Q_{i}$, we calculate
$\delta Q_{m,i}:=\text{negative part of}(Q_{s}+\eta_{0}I+Q_{i}).$
If $\|\lambda_{\max}(\delta Q_{m,i})\|_{L^{\infty}}<\delta_{\eta}:=\eta_{1}-\eta_{0}$, we terminate the iteration.
Otherwise, we approximate $\delta Q_{m,i}$ by a trigonometric polynomial, denoted $\delta Q_{i}$. Specifically, we take $\delta Q_{i}$ to be the Fejér transform of $\delta Q_{m,i}$ in both $r$ and $\theta$, which is equivalent to convolving with the kernel:
@@
Finally, we can obtain a smooth function $Q$ satisfying (6.1).
@@ 6.2 Construction of spectral basis
Since the $Q$-norm $\|\cdot\|_{Q}$ is controlled by the standard $L^{2}$-norm, we begin by considering the following eigenvalue problem in Rayleigh quotient form:
@@
The proof of the lemma is given in Appendix F.
@@
Finally, note that it is necessary to compute all eigenvectors corresponding to eigenvalues smaller than 10. In particular, we need the following estimate to guarantee that all eigenvalues below 10 have been identified: (This lemma is still for the unscaled eigenvalue $\lambda^{s}_{n}$)
@@
    If $l\geqslant 2$ and $\gamma\neq\beta_{l,j}$ for any $j>0$, denote $\tilde{\beta}_{l,j}$ as the sequence obtained by sorting $\{\beta_{l,j},\gamma\}$ in ascending order.
@@
The proof of the lemma is given in Appendix F. In practice, we use the lower bounds of $\alpha_{l,1}^{Y}$ and $\alpha_{l,1}^{Z}$ to determine the largerst index $l$ required, denoted by $l_{\max}$. For every $1\leqslant l\leqslant l_{\max}$, we then search all the intervals $(\beta_{l,j},\beta_{l,j+1})$ and $(\widetilde{\beta}_{l,j},\widetilde{\beta}_{l,j+1})$ for $\alpha_{l,1}^{Z}$ and $\alpha_{l,1}^{Y}$, respectively, until the left endpoint of the interval exceeds the prescribed threshold.
@@ 6.3 Eigenvalue problem
$\langle\psi,Q\psi_{n}^{Q,i}\rangle_{\Omega}=\lambda^{Q,i}_{n}\langle\psi,\psi_{n}^{Q,i}\rangle_{E},\quad\forall\psi\in V_{N}^{i}.$
@@
$\lambda^{U}_{20}\approx 0.456<0.46,\quad\lambda^{v}_{25}\approx 0.453<0.6.$
@@ 6.4 Inversion on the subspace
Finally, we invert the operator $L$ on the subspace, as in the last step in Section 4.2. We numerically solve equation (4.11) by employing the finite element method described in Section 5.5, using the $(\beta,\theta)$ coordinates. We remark that for $\phi_{n}^{v}$, since $\widetilde{\lambda}$ is not explicit in (4.1), we invert instead the operator $P_{\overline{v}}(L^{v}-\overline{\lambda})v$ and leave the small term $\lambda\psi^{v}_{n}$ in the residue $r_{n}^{v}$. Additionally, we must impose the orthogonality constraint:
@@
$\left|-\left(\frac{1}{2}+\overline{\lambda}\right)\phi_{n}^{v}-\frac{1}{2}r\partial_{r}\phi_{n}^{v}+\overline{U}\cdot\nabla_{\operatorname{sph}}\phi_{n}^{v}+\phi_{n}^{v}\cdot\nabla_{\operatorname{sph}}\overline{U}+\nabla_{\operatorname{sph}}q_{n}-\Delta_{\operatorname{sph}}\phi_{n}^{v}-Q\phi_{n}^{v}-\mu\overline{v}\right|,$
@@
When solving for $\phi^{U}_{n}$ with finite element method, we must impose the homogeneous Dirichlet boundary condition, which differs from the condition for $\overline{U}$. Moreover, in the expansion (5.12), the basis functions $\sin((2m-1)\beta)$ must be replaced by $\sin(2m\beta)$. Consequently, the coefficients $\mathrm{B}_{m,l}$-s are required to satisfy a modified condition.
@@ 7 On Rigorous Numerical Verification
In this section we outline the numerical framework we use for validation. The rigorous results of this paper are contained in Sections 16; Section 7 describes the numerical procedure for the ongoing computer-assisted verification. We are implementing these routines and will soon release the code and reproducibility scripts in a public GitHub repository. A subsequent version will report the final constants (we expect only minor adjustments) and the full error budgets from the completed verification.
We focus our effort on verifying the bounds that enter Assumption 1 and the estimates used in Sections 4.2 and 6.26.3. Throughout we work in the spherical coordinates and employ the change of variables
@@
so that all fields are expanded on trigonometric bases in $(\beta,\theta)$. The items below detail how we bound the residuals, $L^{\infty}-$norms, eigenvalue envelopes, and quadrature errors appearing in our verification, in a way that is fully compatible with Assumption 1.
Before we go into the specific details of numerical verification for our method, we would like to make the following observations. First of all, we would like to point out that the numerical verification for the nonuniqueness problem of Leray-Hopf solutions of the 3D Navier–Stokes equations is much simpler than that for the 3D Euler singularity program in [12, 14, 13]. This is largely due to the fact that the 3D Euler singularity problem presents a much more challenging problem. There is no diffusion to provide an additional damping effect, so the stability framework must combine $L^{\infty}$ and $C^{1/2}$ controls with singular weights, to extract a small amount of damping from local terms. However, one has to control multiple nonlocal singular integral operators that do not contribute to stability of the linearized operator. Because local damping is weak, one is forced to obtain very sharp stability constants and use space-time linearized solves to control a finite–rank perturbation of the linearized operator.
In contrast, in the present Navier–Stokes setting, the viscous term yields coercivity, so we can work in the norms regular $L^{2}$ and $H^{1}$ without singular weights. The linearized operator is coercive in $L^{2}$ with a relatively large coercivity coefficient. For $U_{1}$, the coecivity coefficient in the norm $L^{2}$ is given by $1/4-\eta-M_{2}^{U}\approx 0.054$, where we take $\eta=0.191$ and obtain an upper bound for $M_{2}^{U}\leqslant 0.005$. For $v_{1}$, the coercivity coefficient is given by $M_{5}^{v}\approx 0.16$, see Section 4.5 for more details. Our a posteriori bounds close using only the residuals $L^{2}$ of $\overline{U}$ and $\overline{v}$ (plus a few pointwise $L^{\infty}$ envelopes), rather than the weighted $L^{\infty}/C^{1/2}$ controls required in the case of the 3D Euler singularity. The discrete operators we assemble are modest in size (about $1200\times 1200$), and the finite-rank verification reduces to a static generalized eigenproblem of size $25\times 25$. Thanks to strong viscous damping and the effective low-rank approximation of the compact operator $L_{2}^{U}$ using a small number of generalized eigenfunctions $\psi_{j}^{Q,u}$ corresponding to the largest eigenvalues, retaining $25$ modes suffices to close the estimates via the Schauder fixed point theorem.
Moreover, because we approximate $\overline{U}$ and $\overline{v}$ by finite trigonometric tensor products in $(\beta,\theta)$, the residuals $E^{U}$ and $E^{v}$ admit analytic $L^{2}$ evaluation, and many pointwise quantities follow from explicit trigonometric formulas. The only delicate numerical step is the QQ–inner product along the $\beta$–direction under the change of variable $r=\tan\beta$, for which we use a fourth-order composite quadrature together with a rigorous $W^{4,\infty}$ bound on the scalar integrand $(Q\psi_{i}^{Q,u})\cdot(Q\psi_{j}^{Q,u})$ to certify the quadrature error that enters the small matrices. These structural advantages, including viscous coercivity, low-dimensional spectral checks, and analytic trigonometric calculus, make our verification far lighter than that in the 3D Euler singularity case.
@@ 7.1 Analytic $L^{2}$ residuals via a trigonometric representation
We expand the approximate self-similar profile $\overline{U}$ and the (generalized) eigenfunction $\overline{v}$ as finite trigonometric sums
@@
where each coefficient $\mathbf{A}^{\star}_{pq}\in\mathbb{R}^{3}$ collects the three components of $\overline{U}$, and $\mathcal{P}_{\beta},\mathcal{P}_{\theta}\subset\mathbb{N}$ are finite index sets. and similarly for $\overline{v}$. Because sums, products, and derivatives of sines/cosines remain finite trigonometric sums, every term in the residuals $E^{U}$ and $E^{v}$ stays in the same class. The only division introduced arises from the Jacobian in the change of variables $r=\tan\beta$, which produces factors of the form either $\sin(n\beta)/\sin(\beta)$ or $\cos((2m+1)\beta)/\cos(\beta)$. Note that we never have $\cos((2m)\beta)/\cos(\beta)$ in our equations. Such terms will be unbounded for $\theta\approx\pi/2$. For $\sin(n\beta)/\sin(\beta)$, we know that this ratio admits an exact trigonometric expansion, so no true division remains. Indeed, recalling that
@@
In particular, the ratio is a finite cosine polynomial in $\beta$ (no division remaining). Hence every such Jacobian factor can be represented exactly as a finite trigonometric sum and all inner products needed for
$\|E^{U}\|_{L^{2}},\qquad\|E^{v}\|_{L^{2}}$
are evaluated analytically.
We stress again the importance of imposing the correct boundary condition (5.4) for $\widetilde{v}$ at $\theta=0$, If we allow $\widetilde{v}$ to have modes $sin(m\theta)$ or $cos(m\theta)$ with $m\geqslant 2$, then the equation for $\widetilde{v}$ will generate some unbounded terms that would lead to an unbounded residual error $E^{v}$.
@@ 7.2 Certified $L^{\infty}$ bounds on trigonometric polynomials
Let $f$ be any scalar component of $\overline{\phi}_{j}$, $\overline{v}$, or $\overline{U}$. Partition the $\beta$ (or $\theta$) interval into $M$ subintervals $x_{i},x_{i+1}]$ of uniform mesh size $h$, and denote $f_{i}=f(x_{i})$. In each interval $x_{i},x_{i+1}]$, using the classical error bound for linear interpolation with the Lagrange (Cauchy) form of the remainder, we obtain the following.
$|f(x)|\leqslant\max\{f_{i},f_{i+1}\}+\frac{h^{2}}{8}\,\|f^{\prime\prime}\|_{L^{\infty}([x_{i},x_{i+1}])},\qquad x\in[x_{i},x_{i+1}],$
where $h=x_{i+1}-x_{i}$ and $f_{i}=f(x_{i})$.
@@
For a trigonometric polynomial with the highest wavenumber $N$ we have the classical bound $\|f^{\prime\prime}\|_{\infty}\leqslant N^{2}\|f\|_{\infty}$, which gives the local envelope
$|f(x)|\leqslant\frac{\max\{f_{i},f_{i+1}\}}{1-\tfrac{(Nh)^{2}}{8}}\!,\qquad\Rightarrow\quad\|f\|_{\infty}\leqslant\max_{0\leqslant i<M}\frac{\max\{f_{i},f_{i+1}\}}{1-\tfrac{(Nh)^{2}}{8}}.$
@@
Taking $Nh\ll 1$ makes the amplification factor close to $1$. In our runs we use a $300\times 150$ grid in $(\beta,\theta)$, so with fixed spectral cutoff $N$ we pick $M$ large enough that $h=\frac{\pi}{2M}\ll N^{-1}$. In our computation, we take $N=600$ for $\beta$ and $N=300$ for $\theta$, and choose $M=4096$ for $\beta$ and $M=2048$ for $\theta$. All pointwise evaluations $f_{i}$ are computed with interval arithmetic to enclose round-off.
7.3 Bounds for $\|\overline{v}\|_{\infty}$, $\|\nabla\overline{v}\|_{\infty}$, and $\|\nabla\overline{\phi}_{j}\|_{\infty}$
Since $\overline{v}$ and $\overline{\phi}_{j}$ are trigonometric in $(\beta,\theta)$, we evaluate their point values and first derivatives analytically at the grid and then certify global suprema by (7.2). All evaluations carry interval arithmetic so that the reported $L^{\infty}$ bounds are rigorous.
7.4 Far-field decay from the $(r,\beta)$ mapping
The map $r=\tan\beta$ sends $r\to\infty$ as $\beta\uparrow\frac{\pi}{2}$, with
$r=\tan\beta=\frac{1}{\tfrac{\pi}{2}-\beta}+O\!\left(\tfrac{\pi}{2}-\beta\right).$
Through the change of variable $\widehat{U}(\beta,\theta)=\overline{U}/\cos\beta$ and working in the $(\beta,\theta)$ variable, we obtain an approximate solution $\widehat{U}(\beta,\theta)$ as follows:
$\widehat{U}(\beta,\theta)=\sum_{m,n}a_{m,n}\sin((2n{+}1)\beta)\sin(2m\theta).$
Therefor, as $\beta\to\frac{\pi}{2}$ (or $r\to\infty$), we have
$\widehat{U}(\beta,\theta)=A(\theta)+O\!\big((\tfrac{\pi}{2}-\beta)^{2}\big).$
Passing back to the physical self-similar profile uses the smooth factor from the metric,
$\overline{U}(\xi)=\widehat{U}(\beta,\theta)\cos\beta,$
and using the asymptotic property $(\tfrac{\pi}{2}-\beta)^{2}\sim r^{-2}$, we conclude that
$\overline{U}(\xi)=r^{-1}A(\theta)+O(r^{-3}),\qquad r=|\xi|\to\infty.$
The $O(r^{-2})$ term cancels by construction (consistent with transforming back to Cartesian coordinates without generating spurious singular modes). Differentiating $j$ times yields
$|\nabla^{\,j}\overline{U}(\xi)|\lesssim|\xi|^{-j-1},\qquad j\geq 0,$
which matches the far-field hypothesis in Assumption 1. This certifies that the numerically constructed $\overline{U}$ has the required decay property.
7.5 Envelope for the largest eigenvalue of $Q(\beta,\theta)$

### v2

@@ 6 Finite Rank Approximation
$Q_{s}=\frac{\nabla\overline{U}+\nabla^{T}\overline{U}}{2},\quad Q_{s}+\eta_{1}I=Q_{+}-Q_{-},\quad Q_{+}\succeq 0,\quad Q_{-}\succeq 0.$
Here, $Q_{-}$ is the negative part of $Q_{s}+\eta_{1}I$. In previous sections, we choose $Q=Q_{-}$. We truncate the domain $\Omega=[0,R_{0}]\times\left[0,\frac{\pi}{2}\right]$ with $R_{0}=32$ and verify that outside $\Omega$, we have $\lambda_{\max}(Q_{s})<0.199$. Thus, we set $\eta_{1}=0.199$ so that the support of $Q_{-}$ is fully contained in $\Omega$.
Although it is natural to take $Q=Q_{-}$, this choice has the drawback that $Q$ is not smooth, which complicates the future error analysis of numerical integrals. Fortunately, we only need $Q$ to satisfy
@@ 6.1 Construction of $Q$
A naïve approach would be to apply the discrete Fourier transform to $Q_{-}$ and truncate the series. However, this does not ensure $Q_{s}+\eta_{1}I+Q\succeq 0$. To address this, we first select a slightly smaller $\eta_{0}=0.19$ and reconstruct $Q$ using the same idea of a truncation of Fourier series. As long as $L,M$ are sufficiently large, we obtain (6.1).
This method, however, typically requires very large $L,M$, which is undesirable. To improve efficiency, we fix $L=150,M=300$ and use an iterative method. We initialize $Q_{0}=0$ with $\eta_{0}=0.19$. For every iteration $Q_{i}$, we calculate
$\delta Q_{-,i}:=\text{negative part of}\,(Q_{s}+\eta_{0}I+Q_{i}).$
If $\|\lambda_{\max}(\delta Q_{-,i})\|_{L^{\infty}}<\delta_{\eta}:=\eta_{1}-\eta_{0}$, we terminate the iteration.
Otherwise, we approximate $\delta Q_{-,i}$ by a trigonometric polynomial, denoted $\delta Q_{i}$. Specifically, to guarantee that $\delta Q_{i}$ is positive, we take $\delta Q_{i}$ to be the Fejér transform of $\delta Q_{-,i}$ in both $r$ and $\theta$, which is equivalent to convolving with the kernel:
@@
Finally, we can obtain a smooth function $Q$ satisfying (6.1) within 5 iterations.
@@ 6.2 Construction of spectral basis
Since the $Q$-norm $\|\cdot\|_{Q}$ is controlled by the standard $L^{2}$-norm, we begin by considering the following eigenvalue problem in the Rayleigh quotient form:
@@
The proof of the lemma is given in Appendix E.
@@
Finally, note that it is necessary to compute all eigenvectors corresponding to eigenvalues smaller than 10. In particular, we need the following estimate to guarantee that all eigenvalues below 10 have been identified: (This lemma is still for the unscaled eigenvalue $\lambda^{s}_{n}$).
@@
    If $l\geqslant 2$ and $\gamma\neq\beta_{l,j}$ for any $j>0$, denote $\tilde{\beta}_{l,j}$ as the sequence obtained by sorting $\{\beta_{l,j},\gamma\}$ in an ascending order.
@@
The proof of the lemma is given in Appendix E. In practice, we use the lower bounds of $\alpha_{l,1}^{Y}$ and $\alpha_{l,1}^{Z}$ to determine the largerst index $l$ required, denoted by $l_{\max}$. For every $1\leqslant l\leqslant l_{\max}$, we then search all the intervals $(\beta_{l,j},\beta_{l,j+1})$ and $(\widetilde{\beta}_{l,j},\widetilde{\beta}_{l,j+1})$ for $\alpha_{l,1}^{Z}$ and $\alpha_{l,1}^{Y}$, respectively, until the left endpoint of the interval exceeds the prescribed threshold.
@@ 6.3 Eigenvalue problem
$\langle\psi,Q\psi_{n}^{Q,i}\rangle_{\Omega}=\lambda^{Q,i}_{n}\langle\psi,\psi_{n}^{Q,i}\rangle_{E},\quad\forall\psi\in V_{N}^{i},$
@@
$\lambda^{U}_{20}\approx 0.456<0.46,\quad\lambda^{v}_{25}\approx 0.453<0.46.$
@@ 6.4 Inversion on the subspace
Finally, we invert the operator $L$ on the subspace spanned by the finitely many of eigenfunctions, as in the last step in Section 4.2. We numerically solve equation (4.11) by employing the finite element method described in Section 5.5, using the $(\beta,\theta)$ coordinates. We remark that for $\phi_{n}^{v}$, since $\widetilde{\lambda}$ is not explicit in (4.1), we invert instead the operator $P_{\overline{v}}(L^{v}-\overline{\lambda})v$ and leave the small term $\lambda\psi^{v}_{n}$ in the residue $r_{n}^{v}$. Additionally, we must impose the orthogonality constraint:
@@
$\left|-\left(\frac{1}{2}+\overline{\lambda}\right)\phi_{n}^{v}-\frac{1}{2}r\partial_{r}\phi_{n}^{v}+\overline{U}\cdot\nabla_{\operatorname{sph}}\phi_{n}^{v}+\phi_{n}^{v}\cdot\nabla_{\operatorname{sph}}\overline{U}+\nabla_{\operatorname{sph}}q_{n}-\Delta_{\operatorname{sph}}\phi_{n}^{v}-Q\psi_{n}^{v}-\mu\overline{v}\right|,$
@@
When solving for $\phi^{U}_{n}$ with a finite element method, we must impose the homogeneous Dirichlet boundary condition, which differs from the condition for $\overline{U}$. Moreover, in the expansion (5.17), the basis functions $\sin((2m-1)\beta)$ must be replaced by $\sin(2m\beta)$. Consequently, the coefficients $\mathrm{B}_{m,l}$-s are required to satisfy a modified condition.
@@ 7 On Rigorous Numerical Verification
In this section, we outline the framework we use for rigorous computer-assisted validation; in particular, the rigorous justification for the constants in Section 4.5. We release the code and reproducibility scripts in a public GitHub repository. Throughout, we work in the spherical coordinates and employ the change of variables
@@
so that all fields are expanded on trigonometric basis functions in $(\beta,\theta)$. The items below detail how we bound the residuals, $L^{\infty}-$norms, eigenvalue envelopes, and quadrature errors appearing in our verification, in a way that is fully compatible with the verifications in Section 4.5. We remark that verifications concerning the finite-dimensional matrices are much simpler, already sketched in Section 4 and referenced in detail in the codes.
The high-level idea and organization of this section are as follows: we first discuss estimates of numerical integration and, in particular, $L^{2}$-residues in Section 7.1 and 7.2. We leverage the trigonometric representation in the profiles and use exact analytic integration if possible in Section 7.1, and discuss how we perform numerical quadrature with certified error bounds using $L^{\infty}$ bounds on higher-order derivatives in Section 7.2. Then we detail in the remaining part of this section on how to build $L^{\infty}$ estimates. We discuss the classical linear interpolation bound in 2D using higher-order derivatives in Section 7.3 and discuss in particular the use of the Bernstein inequality for trigonometric functions on bounds for these higher-order derivatives. Concerning the construction and the finite-rank approximation of $Q$ and its associated spectral bound estimate, we need to use a coordinate transform on top of the linear interpolation bound in Section 7.4. Finally, for Bessel functions associated with the eigenpairs constructed from the finite-rank approximation, we derive $L^{\infty}$ estimates based on the known spherical Bessel integral representation and plane-wave expansions via Legendre polynomials in Section 7.5.
@@ 7.1 Analytic $L^{2}$ residuals via a trigonometric representation
We expand the approximate self-similar profile $\overline{U}$ and the eigenfunction $\overline{v}$ as finite trigonometric sums
@@
where each coefficient $\mathbf{A}^{\star}_{pq}\in\mathbb{R}^{3}$ collects the three components of $\overline{U}$, and $\mathcal{P}_{\beta},\mathcal{P}_{\theta}\subset\mathbb{N}$ are finite index sets. Similarly, we have an expansion for $\overline{v}$. Because sums, products, and derivatives of sines/cosines remain finite trigonometric sums, every term in the residuals $E^{U}$ and $E^{v}$ stays in the same class. The only division introduced arises from the Jacobian in the change of variables $r=\tan\beta$, which produces factors of the form either $\sin(n\beta)/\sin(\beta)$ or $\cos((2m+1)\beta)/\cos(\beta)$. Note that we never have $\cos((2m)\beta)/\cos(\beta)$ in our equations: such terms will be unbounded for $\beta\approx\pi/2$. For $\sin(n\beta)/\sin(\beta)$, we know that this ratio admits an exact trigonometric expansion, so no true division remains. Indeed, recalling that
@@
In particular, the ratio is a finite cosine polynomial in $\beta$. Hence, every such Jacobian factor can be represented exactly as a finite trigonometric sum, and we can evaluate certain integrals analytically. We remark that while in principle, inner products for $\|E^{U}\|_{L^{2}},\|E^{v}\|_{L^{2}}$ could be evaluated analytically, we choose to use numerical integration for the sake of simplicity; see the next subsection.
Finally we note that we need to keep track of the max frequency. Only products of two trigonometric polynomial increase the max frequency. Therefore, the term with the largest frequency is the transport term $\overline{U}\cdot\nabla\overline{U}$ with max frequency smaller than $2M+10$, where we include a constant wavenumber to account for Jacobians. Here, $M$ stands for the largest frequency of $\overline{U}$. The benefit is that we can easily use the Bernstein inequality to estimate the $L^{\infty}$ bound of any derivative of a trigonometric polynomial, which brings huge convenience for the error estimate of numerical integrals in the following sections.
7.2 Quadrature for numerical integrals and its error
Now we deal with numerical integrals. In contrast to Section 7.1, due to a multiplication of $Q$, the integrals related to the finite rank approximation cannot be evaluated analytically, and we use numerical integrations. For a function $f$, we use the composite Newton–Cotes 7-point rule in each direction, denoted by $\mathrm{Quad}(f)$. A 1D version can be written with an error estimate as
$\int_{a}^{b}f(x)=\frac{h}{140}\Big(41f(x_{0})+216f(x_{1})+27f(x_{2})+272f(x_{3})+27f(x_{4})+216f(x_{5})+41f(x_{6})\Big)-\frac{9}{1400}\,h^{9}\,f^{(8)}(\xi),$
where $\xi\in(a,b),h=\frac{b-a}{6},x_{i}=a+ih\ \ (i=0,\dots,6)$. Let $h_{\beta}$ and $h_{\theta}$ denote the step size in $\beta$ and $\theta$ directions, respectively. We know that the error is controlled by
$\left|\int_{0}^{\frac{\pi}{2}}\int_{0}^{\frac{\pi}{2}}f(\beta,\theta)\mathrm{d}\beta\mathrm{d}\theta-\mathrm{Quad}(f)\right|\leqslant\frac{3\pi^{2}}{11200}(\|\partial_{\beta}^{8}f\|_{L^{\infty}}h_{\beta}^{8}+\|\partial_{\theta}^{8}f\|_{L^{\infty}}h_{\theta}^{8}).$
In many cases, we need to calculate the $L^{2}$ inner product of two functions $f$ and $g$, given by
$\left|\int_{0}^{\frac{\pi}{2}}\int_{0}^{\frac{\pi}{2}}f(\beta,\theta)g(\beta,\theta)\sin\theta\mathrm{d}\beta\mathrm{d}\theta-\mathrm{Quad}(fg\sin\theta)\right|\leqslant\frac{3\pi^{2}}{11200}(F_{\beta}(f,g)h_{\beta}^{8}+F_{\theta}(f,g,\sin\theta)h_{\theta}^{8}),$
where the bound of $\|\partial^{8}_{t}(f_{1}f_{2}\cdots f_{n})\|,t=\beta,\theta$ can be derived via Leibniz’s rule as
$F_{t}(f_{1},f_{2},\ldots,f_{n})\leqslant\sum_{i_{1}+\cdots+i_{n}=8}\left(\begin{array}[]{c}8\\ i_{1},\ldots,i_{n}\end{array}\right)\prod_{j=1}^{n}\|\partial_{t}^{i_{j}}f_{j}\|_{L^{\infty}}.$
In the subsequent subsections, we will detail how to obtain tight $L^{\infty}$ estimates on the functions and their derivatives.
@@ 7.3 Certified $L^{\infty}$ bounds on trigonometric polynomials
Let $f$ be any scalar component of $\bar{\phi}_{j}$, $\bar{v}$, or $\bar{U}$. Partition the $\beta$ (and similarly for $\theta$) interval into $N_{\beta}$ subintervals $\beta_{i},\beta_{i+1}]$ of uniform mesh size $h_{\beta}:=\beta_{i+1}-\beta_{i}$, and denote $f_{i,j}=f(\beta_{i},\theta_{j})$. In each square $I_{i,j}:=[\beta_{i},\beta_{i+1}]\times[\theta_{j},\theta_{j+1}]$, using the classical error bound for linear interpolation with the Lagrange (Cauchy) form of the remainder, we obtain the following estimate:
$|f(\beta,\theta)|\leqslant\max\{f_{i,j},f_{i+1,j},f_{i,j+1},f_{i+1,j+1}\}+\frac{h_{\beta}^{2}\|f_{\beta\beta}\|_{L^{\infty}(I_{i,j})}+h_{\theta}^{2}\|f_{\theta\theta}\|_{L^{\infty}(I_{i,j})}}{8},$ (7.1)
for all $(\beta,\theta)\in I_{i,j}$, where $f_{i,j}=f(\beta_{i},\theta_{j})$.
@@
For a 1D trigonometric polynomial with the highest wavenumber $M$ we have the classical bound via Bernstein’s inequality that $\|f^{\prime\prime}\|_{\infty}\leqslant M^{2}\|f\|_{\infty}$. Taking the supremum and using Bernstein’s inequality, we arrive at
$\|f\|_{\infty}\leqslant(\max_{0\leqslant i\leqslant N_{\beta},0\leqslant j\leqslant N_{\theta}}f_{i,j})/\left(1-\tfrac{(M_{\beta}h_{\beta})^{2}+(M_{\theta}h_{\theta})^{2}}{8}\right).$ (7.2)
Taking $M_{\beta}h_{\beta},M_{\theta}h_{\theta}\ll 1$ makes the amplification factor close to $1$.
To evaluate the upper bound of derivatives of $f$, we can use similar techniques via finite differences. We take $f_{\theta}$ as an example. We first evaluate $f$ on the mesh grids. Then we know that for every $i,j$, there exists a $\xi_{i,j}\in[\theta_{j},\theta_{j+1}]$ such that
$f_{\theta}(\beta_{i},\xi_{i,j})=\frac{f_{i,j+1}-f_{i,j}}{h_{\theta}}.$
Although we do not know the exact value of $\xi_{i,j}$, since $|\xi_{i,j+1}-\xi_{i,j}|<2h_{\theta}$, we can still use the same argument to estimate $\|f_{\theta}\|_{L^{\infty}}$. In fact, as in the previous subsection, we need to obtain sharp estimates of as high as 8th-order derivatives of functions. We can use Bernstein’s inequality to get a crude estimate of the 12th-order derivatives before applying (7.1) to get, recursively, a tighter estimate of 10th-order and 8th-order derivatives.
@@
Recall that in our run, we use a $300\times 150$ grid in $(\beta,\theta)$ for the numerical construction of the approximate solutions. So with a fixed spectral cutoff $M$, we pick $N$ as the finer certification grid $(N_{\beta},N_{\theta})$ used to obtain rigorous $L^{\infty}$ bounds for trigonometric polynomials via the interpolation/Bernstein enclosure, large enough that $h=\frac{\pi}{2N}\ll M^{-1}$. In our computation, we take $M_{\beta}=600$ and $M_{\theta}=300$, and choose $N_{\beta}=4096$ for $\beta$ and $N_{\theta}=2048$ for $\theta$. All pointwise evaluations $f_{i,j}$ are computed with interval arithmetic to enclose round-off.
7.4 Largest eigenvalue estimate: $L^{\infty}$}-estimate via coordinate transform